find the equation for a hyperbola with the vertices (2,0) and (-2,0) with the asymptotes y=7/2x
Answer by lwsshak3(11628)

find the equation for a hyperbola with the vertices (2,0) and (-2,0) with the asymptotes y=7/2x
**
This is a hyperbola with horizontal transverse axis:
Its standard form of equation: (x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of center
For given hyperbola:
center: (0,0)
length of horizontal transverse axis=4 (from -2 to 2)=2a
a=2
a^2=4
given slope of asymptotes for hyperbolas with horizontal transverse axis=b/a=7/2
b=7a/2=7*2/2=14/2=7
b^2=49
equation of given hyperbola:
x^2/4-y^2/49=1